OpenStudy (anonymous):

factor completely: 11x^2-99

OpenStudy (anonymous):

(x+9)(x+11) ?

OpenStudy (kropot72):

First take out the common factor 11. Can you do that?

OpenStudy (zzr0ck3r):

if you foil (x+9)(x+11) you will not get the original, im only sayng this to let you know you can check your answers on these

OpenStudy (anonymous):

k

OpenStudy (zzr0ck3r):

what is 11/11 and 99/11?

OpenStudy (anonymous):

idk than

OpenStudy (anonymous):

1 an 9

OpenStudy (kropot72):

11(x^2 - 9) Can you recognise what the (x^2 - 9) is?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

so (x-1)(x+11)

OpenStudy (zzr0ck3r):

so you said 11(11/11x^2+99/11) what does this simplify too?

OpenStudy (anonymous):

now i'm getting confused

OpenStudy (anonymous):

isn't the anser (x-1)(x-11)

OpenStudy (zzr0ck3r):

simplify what I wrote, when you factor out a number you are deviding the inside terms by that number , and then put that number on the outside. 11(x^2+9)

OpenStudy (kropot72):

(x^2 - 9) is the difference of two squares and factorise to (x + 3)(x - 3) The final answer is: 11(x + 3)(x - 3)

OpenStudy (zepp):

\[11x^2-99\]\[11(x^2)-(11)9\] By using the fact that \[a(b) - c(b) = b(a-c)\] We can do the same thing with this 11 is the common fact, so c \[11(x^2-9)\]

OpenStudy (zzr0ck3r):

no foil your answer out and it wont match

OpenStudy (anonymous):

so is the polynomial prime?

OpenStudy (anonymous):

\[11x^2 -99 = 11(x^2 - 9) = 11(x-3)(x+3)\]

OpenStudy (zepp):

And by using the fact that \[(a-b)(a+b) = a^2 - b^2\] We can factor \[x^2−9 = (x-3)(x+3)\]

OpenStudy (zepp):

So the final answer would be \[11(x-3)(x+3)\]

OpenStudy (anonymous):

is the 11 suppose to stil lb the front to make it 11(x-3)(x+3) ?

OpenStudy (zzr0ck3r):

yes

OpenStudy (zepp):

Yep

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

factor the 11 into the first term and then foil the two binomials and you will get back to where you started

OpenStudy (zzr0ck3r):

err distribute the 11 into the first factor and then foil the two binomails to get the original